OpenStudy (anonymous):

i have log (base x+5) 0.0001= -3 i got x=15 is this correct?

OpenStudy (anonymous):

\[\log_{x+5}(.0001)=-3 \]

OpenStudy (anonymous):

is this how it is written?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ke (x+5) to the right side

OpenStudy (anonymous):

in that case.. (x+5)^-3 = 0.001

OpenStudy (anonymous):

isnt it?

OpenStudy (anonymous):

x+5 = 10

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

0.0001 = (x+5)^-3 \[\sqrt[-3]{0.0001} = x+5\]

OpenStudy (anonymous):

so it would be x=5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

understand it>

OpenStudy (anonymous):

yes in detail better :)

OpenStudy (anonymous):

u want more detail?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

i mean to say \[\sqrt[-3]{0.0001}\neq 1\] then how can you just say x=5 ??? O_o

OpenStudy (anonymous):

its a typo its actually 0.001

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is a typo sorry!

OpenStudy (anonymous):

\[\sqrt[-3]{0.001} = -0.1\]

OpenStudy (anonymous):

=10

OpenStudy (anonymous):

but x=5 was a choice is that not correct his math seemed correct.

OpenStudy (anonymous):

there is not x=10

OpenStudy (anonymous):

x+5=10 so x=5

OpenStudy (anonymous):

no debate

OpenStudy (anonymous):

so if i have log sub 9 27 what would that be 3/2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because 9^(3/2) = 27